# Single-Particle Universality of the Many-Body Spectral Form Factor

### v1

@@ Abstract
We consider systems of fermions evolved by non-interacting unitary circuits with correlated on-site potentials. When these potentials are drawn from the eigenvalue distribution of a circular random matrix ensemble, the spectral form factor (SFF) of the resulting circuit ensemble can be computed exactly without numerical sampling. In the case of the circular unitary ensemble (CUE) we report an exact closed form for the SFF, valid for arbitrary system sizes, and show that it grows through a sequence of exponential ramps. Using exact numerical methods, we find that the circular orthogonal and symplectic ensembles (COE and CSE, respectively) also lead to exponential growth of the SFF. This exponential growth is characteristic of non-interacting systems with random matrix statistics at the single-particle level and, upon introducing interactions, crosses over to a linear ramp consistent with many-body random matrix universality. Our exact results for the SFF provide a baseline for future studies of the crossover between single-particle and many-body random matrix behavior.
Introduction.— Recent advances in experimental physics have ushered in an era of unprecedented control over isolated quantum mechanical systems, embodied in the rise of quantum simulators [1, 2, 3, 4]. In parallel, theorists have devised sophisticated toolsets to identify and characterize different dynamical regimes in many-body systems, often with an eye towards deepening the understanding of fundamental concepts in statistical physics [5, 6, 7]. A cornerstone of these efforts is random matrix theory (RMT), a field whose origins are now decades old [8, 9, 10, 11, 12]. While a number of important conceptual advances have appeared in the intervening years, such as the eigenstate thermalization hypothesis [13, 14, 15, 16, 17, 18], RMT remains a widely-used framework to analyze the spectral properties of quantum systems [18, 19].
Among the popular indicators of random matrix behavior is the spectral form factor (SFF), introduced in pioneering work by Berry and defined as follows [20]. Let $\mathcal{E}$ denote an ensemble of unitary operators; then the SFF of the ensemble is given by $K_{\mathcal{E}}(t)=\langle|\mathrm{Tr}\left(U^{t}\right)|^{2}\rangle_{\mathcal{E}}$, where $U$ is drawn from $\mathcal{E}$ and $\langle\cdot\rangle_{\mathcal{E}}$ denotes ensemble averaging. For interacting systems, the SFF typically decays due to level repulsion down to an $O(1)$ number before entering a regime of linear growth (“the ramp”) beginning at the Thouless time. The ramp persists until the SFF reaches a plateau at the many-body Heisenberg time $t_{H}\sim\mathcal{D}$, where $\mathcal{D}$ is the Hilbert space dimension. Significant effort has gone towards studying form factors in a wide range of physical systems, revealing rich physics as the dynamics flow through single-particle, hydrodynamic, and completely delocalized regimes [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
A challenge in current research is to detect fingerprints of chaotic single-particle dynamics in many-body systems. Early investigations of this dynamical regime were carried out in Refs. [39, 40], which studied the spectral statistics of non-interacting random matrix ensembles embedded in many-body systems. Intriguingly, the form factors of these models grow exponentially in time rather than linearly, with the interpretation that exponential growth is a signature of single-particle chaos in a many-body system. However, while these models are simple in comparison to generic many-body dynamics, computing their form factors has proven difficult and existing methods rely on simplifying properties of the large $\mathcal{D}$ limit.
@@
In this Letter, we introduce many-body random circuit models which describe chaotic single-particle dynamics and whose form factors can be computed exactly without any simplifying approximations or limits. For circuit parameters drawn from the CUE, we report a simple analytic form for the SFF which consists of a sequence of exponential ramps. After discussing this result and its consequences, we conjecture and numerically verify scaling ansatz for other circular ensembles. Finally, we argue that our results can be applied to interacting Hamiltonian and Floquet systems, including a kicked Ising model which we analyze numerically. Our results hence open new possibilities for detecting single-particle chaos in many-body spectra and dynamics.
@@
where $n_{i}$ is a fermion number operator. The many-body quasienergies of $\mathcal{U}$ are determined by the fermion occupations, $\Theta(\bm{n})=\bm{n}\cdot\bm{\theta}$. We are interested in the spectral statistics of this random matrix ensemble, particularly as they are encoded in the SFF,
@@
where the trace is taken over the many-body Hilbert space and $\langle\cdot\rangle_{\mathcal{U}}$ denotes averaging over $\mathcal{U}$, that is, over the $\theta_{j}$. In a minor abuse of notation, we have referred to the ensemble as well as a representative member of the ensemble with the same symbol, $\mathcal{U}$. Note that the SFF is invariant under gauge transformations $\theta_{j}\to\theta_{j}+2\pi m_{j}\;(m_{j}\in\mathbb{Z})$ only if $t$ is an integer, which we assume unless otherwise noted.
Several remarks concerning this model are in order. One might guess that the ensemble-averaged properties of $\mathcal{U}$ are consistent with integrability since the circuits are non-interacting; however, this is incorrect. To see this, note that the spectral properties of the single-fermion sector are characterized by the random matrix statistics of the CUE. In particular, the single particle quasienergies have a joint distribution given by [41, 42]
@@
Both the quasienergy distribution (3) and single-particle SFF (4) are inconsistent with Poisson statistics. While neither of these results speak directly to the many-body spectral statistics of $\mathcal{U}$, the full SFF $K_{\mathcal{U}}(L,t)$ is also inconsistent with integrability. Moreover, it is known [42] that the Gaussian unitary ensemble (GUE) and CUE have the same asymptotic level spacing distributions in the large $L$ limit; hence, while the GUE is often used to model Hamiltonian chaos, one could choose to study the CUE in its place while maintaining many features of the GUE (see Fig. 3 and surrounding discussion).
Exact Spectral Form Factor.— We have evaluated the many-body SFF of Eq. (2) in closed form for parameters drawn from the CUE. At $t=0$, the SFF is given by $4^{L}$, while for any integer $t>0$, we set $N=\lfloor L/t\rfloor$ and find
@@
This result is exact and requires no approximations of any kind. Comparisons of (5) with numerical sampling at fixed $L$ are presented in Fig. 1. Later in this Letter, we will outline some of the key steps leading to (5) and present full details in a separate publication [44]. Here, we analyze the consequences of this result.
Before reaching a plateau $K_{\mathcal{U}}(L,t\geq L)=2^{L}$, the SFF (5) consists of a sequence of exponential ramps with growth rates that depend on $L/t$. To see this, fix $L$ and make a list of its divisors in ascending order: $1=t_{1}<t_{2}<\cdots<t_{M}=L$. Next choose a time $0<t\leq L$ and find the divisors of $L$ which satisfy $t_{j}\leq t\leq t_{j+1}$. Defining $N_{j}=L/t_{j}\in\mathbb{Z}$, Eq. (5) yields
@@
where $\lambda_{j}\equiv(N_{j}+1)\ln(N_{j}+1)-N_{j}\ln(N_{j}+2)$ is the growth rate for $t\in\left[t_{j},t_{j+1}\right)$. In particular, the growth rate jumps when $t$ divides $L$. At late times, meaning $L/t\sim O(1)$, most values of $t$ do not divide $L$ and the SFF is simply described in terms of piecewise exponential functions. At early times, $t/L\ll 1$, the density of divisors of $L$ is large and it is convenient to find an alternative representation for the SFF. When $L/t\in\mathbb{Z}$,
@@
By interpolating this result for times in between divisors of $L$, this form provides a qualitative approximation to the SFF even at late times (see Fig. 1).
@@
This obviously leads to a scaling collapse (see Fig. 1) since our results are exact; we will see nontrivial scaling collapses later (see Figs. 2 and 4).
Calculation Overview.— Here we provide a brief sketch of the calculations which lead to Eq. (5). A full presentation of this and related computations will be published separately [44]. The SFF (2) can be rewritten as
@@
where $P(\bm{\theta})$ is the joint distribution function in Eq. (3). Using well-known properties of the CUE available in texts such as Ref. [42], the SFF admits what we call a moment expansion,
@@
where $\mathtt{R}_{n}$ is the $n$-point single-particle correlation function of the CUE and we refer to $\mathtt{r}_{n}(L,t)$ as the $n$th moment of the SFF.
Thus far, this procedure can be applied with simple modifications to any random matrix ensemble; for example, Ref. [39] implemented a similar analysis for the GUE. However, the CUE has special properties which allow for an exact derivation of (5) for any choice of $L$. In particular, the computation of the moments $\mathtt{r}_{n}(L,t)$ in this case can be mapped onto a combinatorial problem involving the embedding of hardcore particles into a one-dimensional system. Using this combinatorial representation, we have shown that the SFF satisfies the following factorization identity. Given a system size $L$ and a time $t$, write $L=Nt+r$ for some integers $N$ and $r\in\left[0,t-1\right]$; then,
@@
Combined with a proof that $K_{\mathcal{U}}(L,1)=L+1$, this is sufficient to reconstruct the claimed result (5).
Circular orthogonal and symplectic ensembles.— The procedure for generating random non-interacting unitaries for the COE and CSE closely parallels the structure we employed for the CUE. One simply draws a unitary from the appropriate circular ensemble, computes its eigenvalues, and uses them to define $\mathcal{U}$ in Eq. (1). These ensembles exhibit correlated single-particle quasienergies that are linearly related to the many-body quasienergies, so the scaling arguments for the CUE apply and we expect that
@@
for an undetermined function $f$ which depends on the random matrix ensemble under consideration. To study this, we have developed transfer matrix methods which compute the SFF exactly. These methods are derived in detail in our companion paper [44]; here we simply present the results in Fig. 2. Our numerics indicate that the ansatz (13) achieves a scaling collapse for both the COE and CSE models. We leave it to future work to identify the SFF scaling functions.
@@
Comparison with non-interacting GUE fermions.— While our main result (5) holds only for the CUE, it serves as a useful baseline of comparison with other random matrix ensembles, particularly the GUE. Here we will provide a numerical comparison of their SFFs. To study the GUE, we follow the recipe of Ref. [39]. First, draw an $L\times L$ Hermitian matrix $h$ from the GUE and define the many-body fermion Hamiltonian,
@@
Figure 3: Comparison of the many-body CUE and GUE SFFs. The GUE data (solid curves) is averaged over $10^{7}$ disorder realizations and the CUE prediction is taken from (5). The time axis is normalized by the relevant Heisenberg scales, $t_{H}^{\text{(GUE)}}=\frac{\pi}{\sqrt{2}}L$ and $t_{H}^{\text{(CUE)}}=L$. Inset: a magnified view for $t/t_{H}\geq 1/2$ with linear scales.
Fig. 3 shows a comparison between $K_{\text{GUE}}$ and the exact CUE result (5) for a range of system sizes. In both cases we measure time in units of the relevant single-particle Heisenberg time, where $t_{H}^{1,\text{(CUE)}}=L$ and $t_{H}^{1,\text{(GUE)}}=(\pi/\sqrt{2})L$ [45]. At early times, the GUE results exhibit oscillations which relax towards a ramp which is quantitatively close to the exact CUE result (Fig. 3 inset). This behavior persists up to times which are an $O(1)$ fraction of the single-particle Heisenberg time.
In the thermodynamic limit $L\to\infty$, the correlation and cluster functions of the GUE and CUE are known to agree asymptotically [42]. Combined with the numerical results of Fig. 3, this motivates us to conjecture that the CUE and GUE form factors converge in the thermodynamic limit up to an unknown $O(1)$ fraction of the Heisenberg time. Provided that this conjecture holds, it significantly expands the utility of our results for the CUE as a benchmark for other models, such as chaotic many-body systems with weak interactions.
@@
At the non-interacting point $\tau=0$, the KIM has the SFF (5) and grows exponentially in time. For $\tau=1$, Ref. [22] has studied $\mathcal{U}_{\text{KIM}}$ without introducing correlations among the $\theta_{i}$ and argued that it is maximally chaotic. In the strongly interacting regime, correlations among onsite potentials are irrelevant and we expect to see $t$-linear growth of the SFF. More precisely, the SFF with $\tau=1$ is well-described by the many-body COE, for which the SFF is known exactly,
@@
In contrast to our result (5), this SFF grows approximately linearly in $t$ and a scaling collapse is obtained by rescaling $t\to t/t_{H}$ and $K(L,t)\to K(L,t)/t_{H}$.
@@
Acknowledgements.— We acknowledge fruitful discussions with Thomas Scaffidi, Ceren Dağ, Keisuke Fujii, and Kaoru Mizuta. M.O.F. acknowledges support from the Faculty of Science at the University of Victoria through Thomas E. Baker. L.V. acknowledges support from the Slovenian Research and Innovation Agency (ARIS), Research core funding Grants No. P1-0044, N1-0273, J1-50005 and N1-0369, as well as the Consolidator Grant Boundary-101126364 of the European Research Council (ERC). T. N. I. was supported by JST PRESTO Grant No. JPMJPR2112, JSPS KAKENHI Grant No. JP21K13852, and the Boston University CMT visitors program. This research was supported in part by the International Centre for Theoretical Sciences (ICTS) via participation in the program “Stability of Quantum Matter in and out of Equilibrium at Various Scales” (code: ICTS/SQMVS2024/01).

### v2

@@ Abstract
We consider systems of fermions evolved by non-interacting unitary circuits with correlated on-site potentials. When these potentials are drawn from the eigenvalue distribution of a circular random matrix ensemble, the single-particle sector exhibits chaotic dynamics. We study the corresponding many-body spectral statistics and show that the spectral form factor (SFF) can be computed exactly. Due to the absence of interactions the SFF grows exponentially in time, a result which we demonstrate through simple arguments, scaling collapses, and closed-form evaluation of the SFF. We study the role of interactions by numerically analyzing a kicked Ising model and find that the SFF crosses over to a linear growth regime consistent with many-body random matrix universality. Our exact results for the SFF provide a baseline for future studies of the crossover between single-particle and many-body random matrix behavior.
Recent advances in experimental physics have ushered in an era of unprecedented control over isolated quantum mechanical systems [1, 2, 3, 4]. These developments have complemented and renewed theoretical interest in fundamental dynamical concepts such as chaos and ergodicity [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Famously, it is difficult to even rigorously define these terms and researchers often understand them through the lens of random matrix theory (RMT) and spectral statistics [15, 16, 17, 18, 19, 20]. This perspective on many-body dynamics has proven extremely fruitful, generating active research in high-energy physics, quantum information theory, and condensed matter theory [21, 22, 23].
Historically, the insights derived from RMT have proven particularly successful in analyzing several classes of dynamical systems, including integrable models, chaotic single-particle systems (such as billiards), and the late time dynamics of chaotic many-body systems [24, 25, 26, 27]. More recently, new regimes such as systems governed by hydrodynamics have been studied from the perspective of spectral statistics, particularly as they are encoded in the spectral form factor (SFF) [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. It remains an important problem in many-body dynamics to search for models which are amenable to exact solutions while exhibiting universal dynamical properties.
@@
In this Letter, we analyze models of non-interacting fermions and argue that their many-body statistics can be studied exactly. The exactness of our results makes these models strong candidates to analyze the onset of many-body random matrix universality in the presence of interactions; rigorously analyzing the non-interacting regime is an essential first step in this direction. To that end, we compute the SFFs of our fermion models using exact numerical and analytical approaches. For circuit parameters drawn from the circular unitary ensemble (CUE), we report a simple analytic form for the SFF at integer times which consists of a sequence of exponential ramps. After discussing this result and its consequences, we conjecture and numerically verify scaling ansatz for the circular orthogonal and symplectic ensembles (COE and CSE, respectively). Finally, we argue that our results can be applied to interacting Hamiltonian and Floquet systems, including a kicked Ising model which we analyze numerically. Our results hence open new possibilities for detecting single-particle chaos in many-body spectra and dynamics.
@@
where $n_{i}$ is a fermion number operator. The many-body quasienergies of $\mathcal{U}$ are determined by the fermion occupations, $\Theta(\bm{n})=\bm{n}\cdot\bm{\theta}$. We are interested in the spectral statistics of this random matrix ensemble, particularly as they are encoded in the SFF [51]
@@
where the trace is taken over the many-body Hilbert space and $\langle\cdot\rangle_{\mathcal{U}}$ denotes averaging over $\mathcal{U}$, that is, over $\theta_{j}$. In a minor abuse of notation, we have referred to the ensemble as well as a representative member of the ensemble with the same symbol, $\mathcal{U}$. Throughout this Letter we assume that $t\in\mathbb{Z}$, which guarantees that the SFF is invariant under “gauge transformations” $\theta_{j}\to\theta_{j}+2\pi m_{j},\;m_{j}\in\mathbb{Z}$. However, our methods for computing the SFF can be extended to arbitrary real times as discussed in Ref. [52].
Several remarks concerning this model are in order. One might guess that the ensemble-averaged properties of $\mathcal{U}$ are consistent with integrability since the circuits are non-interacting; however, this is incorrect. To see this, note that the single-fermion sector is characterized by the random matrix statistics of the CUE. In particular, the single-particle quasienergies have the joint distribution [53, 54]
@@
Both the quasienergy distribution (3) and single-particle SFF (4) are inconsistent with Poisson statistics. While neither of these results speak directly to the many-body spectral statistics of $\mathcal{U}$, we will show that the SFF $K_{\mathcal{U}}(L,t)$ is also inconsistent with integrability. Moreover, it is known [54] that the Gaussian unitary ensemble (GUE) and CUE have the same asymptotic level spacing distributions in the large $L$ limit; hence, while the GUE is often used to model Hamiltonian chaos, one could choose to study the CUE in its place while maintaining many features of the GUE (see Fig. 3 and surrounding discussion).
Exact Spectral Form Factor.— We have evaluated the many-body SFF (2) in closed form for parameters drawn from the CUE. At $t=0$, the SFF is given by $4^{L}$, while for any integer $t>0$, we set $N=\lfloor L/t\rfloor$ and find
@@
This result is exact and requires no approximations. Comparisons of (5) with numerical sampling at fixed $L$ are presented in Fig. 1. Later in this Letter, we will outline some of the key steps leading to (5) and present full details in a separate publication [52]. Here, we analyze the consequences of this result.
Before reaching the plateau $K_{\mathcal{U}}(L,t\geq L)=2^{L}$, the SFF (5) consists of a sequence of exponential ramps with growth rates that depend on $L/t$. To see this, fix $L$ and make a list of its divisors: $1=t_{1}<t_{2}<\cdots<t_{M}=L$. Next fix $0<t\leq L$ and find the divisors of $L$ which satisfy $t_{j}\leq t\leq t_{j+1}$. Defining $N_{j}=L/t_{j}\in\mathbb{Z}$, Eq. (5) yields
@@
where $\lambda_{j}\equiv(N_{j}+1)\ln(N_{j}+1)-N_{j}\ln(N_{j}+2)$ is the growth rate for $t\in\left[t_{j},t_{j+1}\right)$. In particular, the growth rate jumps when $t$ divides $L$. At late times,$t/L\sim O(1)$, most values of $t$ do not divide $L$ and the SFF is simply described by piecewise exponential functions. When $t/L\ll 1$, the density of divisors of $L$ is large and it is convenient to find an alternative representation for the SFF. When $L/t\in\mathbb{Z}$,
@@
By interpolating this result for times between divisors of $L$, this form qualitatively approximates the SFF even at late times (see Fig. 1).
@@
This obviously leads to a scaling collapse since our results are exact (see Fig. 1); we will see nontrivial scaling collapses later (see Figs. 2 and 4).
Calculation Overview.— Here we provide a sketch of the calculations which lead to (5). A full presentation of this and related results will be published separately [52]. The SFF (2) can be written as
@@
where $P(\bm{\theta})$ is defined in (3). Using well-known properties of the CUE [54], the SFF admits what we call a moment expansion,
@@
where $\mathtt{R}_{n}$ is the $n$-point single-particle correlation function of the CUE and $\mathtt{r}_{n}(L,t)$ is the $n$th moment.
Thus far, this procedure can be applied with simple modifications to any random matrix ensemble; for example, Ref. [50] implemented a similar analysis for the GUE. However, the CUE has special properties which allow for an exact derivation of (5) for any choice of $L$. In particular, the computation of the moments can be mapped onto a combinatorial problem of hardcore particles in a one-dimensional system. Using this representation, we have shown that the SFF satisfies the following factorization identity. Given a system size $L$ and a time $t$, write $L=Nt+r$ for some integers $N$ and $r\in\left[0,t-1\right]$. Then
@@
Combined with a proof that $K_{\mathcal{U}}(L,1)=L+1$, this implies the claimed result (5).
Other circular ensembles.— The procedure for generating random unitaries for the COE and CSE closely parallels the structure we employed for the CUE. One simply draws a unitary from the appropriate circular ensemble, computes its eigenvalues, and uses them to define $\mathcal{U}$ in Eq. (1). These ensembles exhibit correlated single-particle quasienergies that are linearly related to the many-body quasienergies, so our scaling arguments apply and we expect that
@@
for an undetermined function $f$ which depends on the random matrix ensemble. To study this, we have developed transfer matrix methods which compute the SFF exactly. These methods are derived in detail in Ref. [52]; here we present numerical results in Fig. 2. Our numerics indicate that the ansatz (13) achieves a scaling collapse for both the COE and CSE models. We leave it to future work to identify the scaling functions.
@@
Comparison with non-interacting GUE fermions.— While our main result (5) holds only for the CUE, it serves as a useful baseline of comparison with other random matrix ensembles, particularly the GUE. Here we will provide a numerical comparison of their SFFs. To study the GUE, we follow the recipe of Ref. [50]. First, draw an $L\times L$ Hermitian matrix $h$ from the GUE and define the many-body fermion Hamiltonian
@@
Figure 3: Comparison of the many-body CUE and GUE SFFs. The GUE data (solid curves) is averaged over $10^{7}$ disorder realizations and the CUE prediction is taken from (5). The time axis is normalized by the single-particle Heisenberg scales, $t_{H}^{\text{1, (GUE)}}=\frac{\pi}{\sqrt{2}}L$ and $t_{H}^{\text{1, (CUE)}}=L$. Inset: a magnified view for $t/t_{H}^{(1)}\geq 1/2$ with linear scales.
Fig. 3 shows a comparison between $K_{\text{GUE}}$ and the exact CUE result (5) for a range of system sizes. In both cases we measure time in units of the relevant single-particle Heisenberg time, where $t_{H}^{1,\text{(CUE)}}=L$ and $t_{H}^{1,\text{(GUE)}}=(\pi/\sqrt{2})L$ [56]. At early times, the GUE results exhibit oscillations which relax towards a ramp that is quantitatively close to the exact CUE result (Fig. 3 inset). This behavior persists up to times which are an $O(1)$ fraction of the single-particle Heisenberg time.
In the thermodynamic limit $L\to\infty$, the correlation and cluster functions of the GUE and CUE are known to agree asymptotically [54]. Combined with the numerical results of Fig. 3, this motivates us to conjecture that the GUE form factors follows our exact result for the CUE in the thermodynamic limit up to an unknown $O(1)$ fraction of the many-body Heisenberg time. Provided that this conjecture holds, it significantly expands the utility of our results for the CUE as a benchmark for other models, such as chaotic many-body systems with weak interactions.
@@
At the non-interacting point $\tau=0$, the KIM has the SFF (5) and grows exponentially in time. For $\tau=1$, Ref. [29] has studied $\mathcal{U}_{\text{KIM}}$ without introducing correlations among the $\theta_{i}$ and argued that it is maximally chaotic. In the strongly interacting regime, correlations among $\theta_{j}$ are irrelevant and we expect $t$-linear growth of the SFF. More precisely, the SFF with $\tau=1$ is described by the many-body COE, for which the SFF is known,
@@
In contrast to our result (5), this SFF grows linearly in $t$ and a scaling collapse is obtained by rescaling $t\to t/t_{H}$ and $K(L,t)\to K(L,t)/t_{H}$.
@@
Acknowledgements.
We acknowledge fruitful discussions with Thomas Scaffidi, Ceren Dağ, Keisuke Fujii, and Kaoru Mizuta. M.O.F. acknowledges support from the Faculty of Science at the University of Victoria through Thomas E. Baker. L.V. acknowledges support from the Slovenian Research and Innovation Agency (ARIS), Research core funding Grants No. P1-0044, N1-0273, J1-50005 and N1-0369, as well as the Consolidator Grant Boundary-101126364 of the European Research Council (ERC). T. N. I. was supported by JST PRESTO Grant No. JPMJPR2112, JSPS KAKENHI Grant No. JP21K13852, and the Boston University CMT visitors program. This research was supported in part by the International Centre for Theoretical Sciences (ICTS) via participation in the program “Stability of Quantum Matter in and out of Equilibrium at Various Scales” (code: ICTS/SQMVS2024/01).